OpenStudy (anonymous):

simplify the expression 6n2-5n2+7n2 i know it equals 8n2 but what would this expression be in a simplified form ?

Nnesha (nnesha):

ok so they all have same varialbe and same exponent

Nnesha (nnesha):

variable **

Nnesha (nnesha):

so just combine like terms

OpenStudy (anonymous):

i guess so ? i just put 6-5=1+7=8

Nnesha (nnesha):

6 -7 + 5 = ??

Nnesha (nnesha):

yes so that is your answer \[8n ^{2}\]

sammixboo (sammixboo):

I believed that is already in simplified form, 8n2 Are they exponents?!

OpenStudy (anonymous):

thats in simplified form

OpenStudy (anonymous):

i feel stupid

Nnesha (nnesha):

nope you arenot

sammixboo (sammixboo):

If they are exponents, then \(\rm 8n^2\) is the simplified form And no, you are not stupid :)

sammixboo (sammixboo):

Oh, and... \(\bf \Large \color{magenta}{W}\color{aqua}{e}\color{magenta}{l}\color{aqua}{c}\color{magenta}{o}\color{aqua}{m}\color{magenta}{e}\) \(\bf \large to\) \(\bf \Large \color{Aqua}{O}\color{Aqua}{p}\color{aqua}{e}\color{aqua}{n}\color{magenta}{S}\color{magenta}{t}\color{magenta}{u}\color{magenta}{d}\color{magenta}{y!}\)

Nnesha (nnesha):

lol

Nnesha (nnesha):

sammi you are really fast when its about to latex

sammixboo (sammixboo):

Ah I have an OpenStudy extension called 'OpenStudy Notebook'. I can save all my LaTeX and all their :)

sammixboo (sammixboo):

there* T-T...

Nnesha (nnesha):

0.o smart gurl

sammixboo (sammixboo):

Yes :P There is also a new OpenStudy extension called OpenStudy Quoting. You can instantly quote other users replies and can do some other things :) Check it out soon!

Nnesha (nnesha):

i will love to do i will talk you tomorrow about it ok :)

Nnesha (nnesha):

@Prone632 sorry :(

sammixboo (sammixboo):

See the Quoting extentsion you download shows the features 'Quote' and 'Copy' under any comment in post. Click 'Quote' under a commment and this will pop up.. \(\color{blue}{\text{Originally Posted by}}\) @Nnesha @Prone632 sorry :( \(\color{blue}{\text{End of Quote}}\) If you copy a text it will automatically copy a text into your comment box, and it even works with LaTeX!

Nnesha (nnesha):

\(\color{purple}{AWESOME}\) :) thanks